OpenStudy (anonymous):

what is the binomial probability of the following: P(X < 5) P(X less than or equal to 10) P(4 < X greater than or equal to 9)

OpenStudy (john_es):

Do you mean in a coin experiment? I think the probabily of success must be given.

OpenStudy (anonymous):

Oh yea im sorry. n=12 p=.3 so q=.7

OpenStudy (john_es):

Apply the binomial distribution function, \[P(X<5)=\sum_{k=1}^{4}\binom {12}k p^kq^{12-k}\] For the second part, \[P(X\leq 10)=1-\sum_{k=1}^{2}\binom {12}kp^kq^{n-k}\] For the third part \[P(X\geq 4)=1-\sum_{k=1}^{3}\binom {12}kp^kq^{n-k}\]

OpenStudy (john_es):

Sorry, the sums must be from k=0.

OpenStudy (john_es):

Results, 0.723655 0.747185 0.507484

OpenStudy (anonymous):

the second two arent right

OpenStudy (john_es):

True, it must be, for the second 0.999985, and the third it must be, this one, \[P(4<X\leq 9)=P(X=5)+P(X=6)+P(X=7)+P(X=8)+P(X=9)=0.276138\]